OpenStudy (anonymous):

if the quatity of heat requried to raise the temperature of 70g of a gas from 25 C to 35C at constant pressure is 725J the gas is ?

OpenStudy (abb0t):

I'm misreading this question, what are you looking for? Lol

OpenStudy (anonymous):

I am Looking the GAS Used

OpenStudy (anonymous):

(a) H2 (B) N2 (C) O2 (D) Cl2

OpenStudy (anonymous):

It means We have to Find the Molar Mass of the Gas used Here

OpenStudy (anonymous):

@electrokid @.Sam. @Callisto

OpenStudy (abb0t):

since pressure it constant, it's meaning that you don't need to worry about it.

OpenStudy (abb0t):

Pressure and volume are inversely proportional.

OpenStudy (abb0t):

Try and focus on the relationship of the info given. Does that make sense?

OpenStudy (abb0t):

What formula do you know that can relate that...

OpenStudy (anonymous):

Q=mc (T2-T1)

OpenStudy (anonymous):

i can Only See This one..

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (anonymous):

not the molar mass dude, u gotta find its specific heat. http://www.engineeringtoolbox.com/spesific-heat-capacity-gases-d_159.html and u may use this to find out your gas

OpenStudy (abb0t):

Sorry for the late response, I was working on something last night. But yes, That's the formula you want to start.

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

@chmvijay

OpenStudy (anonymous):

what value did you get for "c"?

OpenStudy (anonymous):

Q = mc dT 725 = 70 *c * 10 C = 725/700

OpenStudy (anonymous):

remember that since this is a gas at const. pressure you will get "C_p"

OpenStudy (anonymous):

nice.. divide that with "R"

OpenStudy (anonymous):

Why Divide by R ?

OpenStudy (anonymous):

(make sure of the units) you can either use a table to see which gas has that "C_p" value or if you divide by "R" you will get a hint as to if it'd be a monoatomic or diatomic gas

OpenStudy (anonymous):

C/R = 0.12

OpenStudy (anonymous):

did you make sure you used the proper units?

OpenStudy (anonymous):

the ratio should be >2

OpenStudy (anonymous):

it is N2

OpenStudy (chmvijay):

:(

OpenStudy (anonymous):

cp values for the given options: (J/gK) a) 14.3 b) 1.04 c) 0.92 d) ?? (dont know) source -> wiki

OpenStudy (chmvijay):

LOL:) with wiki we should not tell this y solving we should tell LOL

OpenStudy (anonymous):

Using the values you quoted: Q = mass * (temperature change) 725 = 70 * C * 10 C = 725/700 = 1.04 Joule per gram per Kelvin cp values for the given options: (J/gK) a) 14.3 b) 1.04 c) 0.92 d) so the measured heat capacity (at constant pressure) matches choice b most closely, making N2 the correct answer.

OpenStudy (anonymous):

I am writing off the top of my bird-brain, so correct me if i am wrong. Let's pretend these gases are ideal gases. so the "constant pressure" should not be much higher than 1 atm, the temperature should not be too low, yada yada. For a diatomic gas, there are 6 degrees of freedom, so the heat capacity (*at constant volume*) or C_v should be 6*1/2 R Joules per mol per Kelvin. why six degrees of freedom? roughly speaking, because there are 3 degrees of freedom for translation, 2 for rotation, 1 for vibration. for an ideal gas, the molar heat capacity at constant *Pressure* is heat capacity at constant *Volume* + R so C_p = 4 R Joule per Kelvin per *mole* Now the specific heat is in units of Joule per Kelvin per *gram*, so we need to convert the unit: specific heat = C_p / (molecular weight) and we see that the unit comes out to be correct. plugging in the values for the examples you gave, we can estimate the specific heats: H2: 17 (versus the measured value of 14.3) N2: 1.2 (1.04) O2: 1.0 (0.92) Cl2: 0.47 the unit is in Joules per Kelvin per gram. i put the measured value in bracket. The estimated values are a bit off, but they show why the specific heat decreases as the molar mass increases.